# Unconditional Density Bounds for Quadratic Norm-Form Energies via Lorentzian Spectral Weights Persistent Heuristics: Mini Monograph I

### 03 Mar 2026 v2

@@ Appendix B Zeros of $L(s,\chi_{5})$ to 70 Decimal Places (Conductor 5, $d=5$)
11 6.64845334472771471612327845997931784729858542324449837237471294671536936.6484533447277147161232784599793178472985854232444983723747129467153693 $3.65\times 10^{-400}$
22 9.83144443288666961634832134745844382188813289401550743042668652269440359.8314444328866696163483213474584438218881328940155074304266865226944035 $5.71\times 10^{-379}$
33 11.958845626083514530265658688262841817293112765723287717381364444659796711.9588456260835145302656586882628418172931127657232877173813644446597967 $2.45\times 10^{-377}$
44 16.033821128384235674593253782248083508749303711393763454343960473652264016.0338211283842356745932537822480835087493037113937634543439604736522640 $1.67\times 10^{-374}$
55 17.566994292325555202701595268144486403485240895508468763100055336820239617.5669942923255552027015952681444864034852408955084687631000553368202396 $6.87\times 10^{-379}$
66 19.540732622784750250378690022998553485526493445007156424107667950060473619.5407326227847502503786900229985534855264934450071564241076679500604736 $4.30\times 10^{-381}$
77 22.227405454459410911877624963081473365280642640560187108353392333259926322.2274054544594109118776249630814733652806426405601871083533923332599263 $1.80\times 10^{-378}$
88 24.588466217408195207656269976081979066247857188503256974763128082376573724.5884662174081952076562699760819790662478571885032569747631280823765737 $3.05\times 10^{-376}$
99 26.776095948004140116523574965270948586289370096100720747141444091727181626.7760959480041401165235749652709485862893700961007207471414440917271816 $4.82\times 10^{-383}$
$10$ 28.461035100177522475186978272325256271200734674528650404495427630566281928.4610351001775224751869782723252562712007346745286504044954276305662819 $3.67\times 10^{-379}$
$11$ 29.707909350480965569230986518657313605970196432418479892145402797880542429.7079093504809655692309865186573136059701964324184798921454027978805424 $5.77\times 10^{-383}$
$12$ 33.000456006870514367949759177211937317376500804631238343447920775517672933.0004560068705143679497591772119373173765008046312383434479207755176729 $2.60\times 10^{-381}$
$13$ 34.728812978904808674143729833981379136564007992078974829862022029197196934.7288129789048086741437298339813791365640079920789748298620220291971969 $7.03\times 10^{-380}$
$14$ 35.868638371812274594595048638876426333384181725562636611431705494857515935.8686383718122745945950486388764263333841817255626366114317054948575159 $8.26\times 10^{-388}$
$15$ 38.129184721436531850151418270372918926216187170437541987827769860251830338.1291847214365318501514182703729189262161871704375419878277698602518303 $4.26\times 10^{-378}$
$16$ 39.560572946403181705055095059954798425566825204646381762421875276690628839.5605729464031817050550950599547984255668252046463817624218752766906288 $1.04\times 10^{-377}$
$17$ 41.842438545791694308509306885312933146785226936948497507540546923147676541.8424385457916943085093068853129331467852269369484975075405469231476765 $1.47\times 10^{-382}$
$18$ 44.031290061441695044700908058424922514756441978971928142545909588119921544.0312900614416950447009080584249225147564419789719281425459095881199215 $3.53\times 10^{-371}$
1919 45.427300082782289388841561909609720649965874840453313174253112295195133545.4273000827822893888415619096097206499658748404533131742531122951951335 $1.04\times 10^{-383}$
2020 46.492727159491405345339199351675833673263826528873731752681947727399791446.4927271594914053453391993516758336732638265288737317526819477273997914 $1.44\times 10^{-379}$
2121 48.345661821067846176548201304493486929945215885626654705310956748852704448.3456618210678461765482013044934869299452158856266547053109567488527044 $7.47\times 10^{-384}$
$22$ 51.087751926746491355258257204131245513931822764598874889643028952539913351.0877519267464913552582572041312455139318227645988748896430289525399133 $4.99\times 10^{-375}$
$23$ 52.125902231316974188604130695930300524969293830813142026749809937021618752.1259022313169741886041306959303005249692938308131420267498099370216187 $1.86\times 10^{-381}$
$24$ 53.830445195442163351054269022338568937339170713109887047091903708283035253.8304451954421633510542690223385689373391707131098870470919037082830352 $6.33\times 10^{-386}$
$25$ 55.589280335404810150964588998492970666633756779445931871191432982474752955.5892803354048101509645889984929706666337567794459318711914329824747529 $6.76\times 10^{-379}$
$26$ 56.838865942890988462381701087979625873787502823421640673322572691726700956.8388659428909884623817010879796258737875028234216406733225726917267009 $3.20\times 10^{-377}$
$27$ 58.386117485358380826793925742317062296785252761744412031426483942467746058.3861174853583808267939257423170622967852527617444120314264839424677460 $3.82\times 10^{-376}$
$28$ 61.138864751921745109095463915572425456187565195524483622740465340768912261.1388647519217451090954639155724254561875651955244836227404653407689122 $2.98\times 10^{-378}$
$29$ 62.132904720233877934071529440383543136671977562274208379683708348891512562.1329047202338779340715294403835431366719775622742083796837083488915125 $3.12\times 10^{-386}$
$30$ 63.709543060981146771798494836335680246320729913038089464768398907875187863.7095430609811467717984948363356802463207299130380894647683989078751878 $6.60\times 10^{-377}$
$31$ 64.637359665374574101088589957514469169614418340871541926806467731667825864.6373596653745741010885899575144691696144183408715419268064677316678258 $3.08\times 10^{-374}$
$32$ 66.747641845937807830433496332522277799056278904231710999500592219175209366.7476418459378078304334963325222777990562789042317109995005922191752093 $1.44\times 10^{-384}$
$33$ 68.590727924825006540833515068030970994309372068387999238433409618374964168.5907279248250065408335150680309709943093720683879992384334096183749641 $7.49\times 10^{-385}$
$34$ 70.115837159193504702254162003288501122534880816675087043045923668594310970.1158371591935047022541620032885011225348808166750870430459236685943109 $1.94\times 10^{-376}$
$35$ 71.605959479243017199477227381292487650674837775870501191061771430349257871.6059594792430171994772273812924876506748377758705011910617714303492578 $7.09\times 10^{-376}$
$36$ 73.350105431468259324986481309687463592380080206128610550028406269746719673.3501054314682593249864813096874635923800802061286105500284062697467196 $2.28\times 10^{-381}$
$37$ 74.293688438016568417563919573323746752656162649388588394478627746168005574.2936884380165684175639195733237467526561626493885883944786277461680055 $9.97\times 10^{-379}$
$38$ 75.534904117228113405345989547687724838297800656054992950178626192713539375.5349041172281134053459895476877248382978006560549929501786261927135393 $1.36\times 10^{-378}$
$39$ 78.072170864382337194051431823672802856073274377354409700849470519410328178.0721708643823371940514318236728028560732743773544097008494705194103281 $2.97\times 10^{-372}$
$40$ 79.786062490097939106371571243171617631261869341105718896147463944735173779.7860624900979391063715712431716176312618693411057188961474639447351737 $3.45\times 10^{-378}$
$41$ 80.567300856410698261373757677661867641972294500859119722736169220177608280.5673008564106982613737576776618676419722945008591197227361692201776082 $3.03\times 10^{-379}$
$42$ 81.923380893154115311967989553905914195519790752160387751303590547918086981.9233808931541153119679895539059141955197907521603877513035905479180869 $5.41\times 10^{-383}$
$43$ 83.699157001874113162883288392240055676480712410299839420963639633825894883.6991570018741131628832883922400556764807124102998394209636396338258948 $3.38\times 10^{-377}$
$44$ 84.952507103754973661846904660485113866224814086703961086051329413504388084.9525071037549736618469046604851138662248140867039610860513294135043880 $8.50\times 10^{-372}$
$45$ 86.809996679338770496962453824123865539107306122702912569153552151872514786.8099966793387704969624538241238655391073061227029125691535521518725147 $9.72\times 10^{-386}$
$46$ 88.281776599709835593621908069365568911371943199556447973711012167487531688.2817765997098355936219080693655689113719431995564479737110121674875316 $7.96\times 10^{-391}$
$47$ 90.376573182128156802963266298552786636345751535570441691904969596990152990.3765731821281568029632662985527866363457515355704416919049695969901529 $2.68\times 10^{-371}$
$48$ 90.715525933146974977269052021613460765711009835158005422179562303740599890.7155259331469749772690520216134607657110098351580054221795623037405998 $4.10\times 10^{-379}$
$49$ 92.450243253774403222656651609940166532625651637342691786476584587897979992.4502432537744032226566516099401665326256516373426917864765845878979799 $9.68\times 10^{-380}$
$50$ 93.394423937828017400982646092448145456052891559124280035714176687016475693.3944239378280174009826460924481454560528915591242800357141766870164756 $5.09\times 10^{-377}$
$51$ 95.970001646037706939832706231680914850344362198662140910938755275316429995.9700016460377069398327062316809148503443621986621409109387552753164299 $3.46\times 10^{-382}$
$52$ 97.331588252898071590582881239627993663344456879633576511321148041911903697.3315882528980715905828812396279936633444568796335765113211480419119036 $1.34\times 10^{-379}$
$53$ 98.227457869901077077061547196562023327793547734165391303478765895425843798.2274578699010770770615471965620233277935477341653913034787658954258437 $7.80\times 10^{-377}$
$54$ 99.838050807536549892458178740443727323094718661873207723162307044585847899.8380508075365498924581787404437273230947186618732077231623070445858478 $2.09\times 10^{-387}$
$55$ 101.1877170969186106879893673648615779856303956787129404948013652775331402101.1877170969186106879893673648615779856303956787129404948013652775331402 $8.26\times 10^{-389}$
$56$ 102.6052879410500105919016338097244199966334774046924761659561257554903097102.6052879410500105919016338097244199966334774046924761659561257554903097 $7.88\times 10^{-387}$
$57$ 103.6661310797677629747891494548931064951051111049995242765910102899167622103.6661310797677629747891494548931064951051111049995242765910102899167622 $1.11\times 10^{-399}$
$58$ 105.9731937977587861297183750394317655839089623673242427608883134019122134105.9731937977587861297183750394317655839089623673242427608883134019122134 $9.41\times 10^{-387}$
$59$ 107.2971036375259575240109457236330672603545936270064737698887145263170400107.2971036375259575240109457236330672603545936270064737698887145263170400 $4.62\times 10^{-386}$
$60$ 108.4981066242615983601377772869913646076568520675142124199793643754286051108.4981066242615983601377772869913646076568520675142124199793643754286051 $1.63\times 10^{-382}$
$61$ 109.5502586071841272439535415119528096196195061162254833594241784691252750109.5502586071841272439535415119528096196195061162254833594241784691252750 $9.23\times 10^{-388}$
$62$ 110.6361976726278518372473829197474157598718943493419745206490444961344614110.6361976726278518372473829197474157598718943493419745206490444961344614 $1.65\times 10^{-384}$
$63$ 112.6157699450494525603420477631256573284904001732412552640056168054705385112.6157699450494525603420477631256573284904001732412552640056168054705385 $1.47\times 10^{-378}$
$64$ 114.2306398723819524205236259115864715634752773232857784617817436415063754114.2306398723819524205236259115864715634752773232857784617817436415063754 $8.49\times 10^{-385}$
$65$ 115.5784139823770731492864674907752976644587354074111098082626495892785777115.5784139823770731492864674907752976644587354074111098082626495892785777 $2.08\times 10^{-382}$
$66$ 116.6356717597837992860716616244380290420850400876742741600872550143161217116.6356717597837992860716616244380290420850400876742741600872550143161217 $5.60\times 10^{-393}$
$67$ 118.3711623063552621519304519550474639914856549056003421824775094839317423118.3711623063552621519304519550474639914856549056003421824775094839317423 $1.09\times 10^{-367}$
$68$ 119.5614916258800030014640284654174824520291754354058843504968760508448926119.5614916258800030014640284654174824520291754354058843504968760508448926 $5.21\times 10^{-386}$
$69$ 120.2817341918527579929633987893433395906991345403172727276829320830538688120.2817341918527579929633987893433395906991345403172727276829320830538688 $6.70\times 10^{-376}$
$70$ 122.0616322838941436366460069970312459167329049609093140518854982869671603122.0616322838941436366460069970312459167329049609093140518854982869671603 $2.75\times 10^{-383}$
$71$ 124.2428371489503574456280462431539381366392074620427334384975693951880954124.2428371489503574456280462431539381366392074620427334384975693951880954 $2.98\times 10^{-375}$
$72$ 125.2815361690544786849959629329922579004633315004421413848554281307878485125.2815361690544786849959629329922579004633315004421413848554281307878485 $4.25\times 10^{-376}$
$73$ 126.3986993784245731137991303757675004132192700408244704238571151980971242126.3986993784245731137991303757675004132192700408244704238571151980971242 $3.04\times 10^{-387}$
$74$ 127.3206348281152351166040431083884153111288443677405189033098037679391082127.3206348281152351166040431083884153111288443677405189033098037679391082 $2.80\times 10^{-381}$
$75$ 129.0483000943208951329053997018161493807340037666172892353902573266257761129.0483000943208951329053997018161493807340037666172892353902573266257761 $1.49\times 10^{-385}$
$76$ 130.1970382668577478738346244894136617196218666973159138428989678080490833130.1970382668577478738346244894136617196218666973159138428989678080490833 $4.04\times 10^{-397}$
$77$ 132.0860053031175217593818811988220993124226919197273881005207706354373250132.0860053031175217593818811988220993124226919197273881005207706354373250 $1.69\times 10^{-374}$
$78$ 133.0302283883319444007316023118379973226202504376344453082369748548055878133.0302283883319444007316023118379973226202504376344453082369748548055878 $5.31\times 10^{-377}$
$79$ 134.9249638941115606306389354461440563212716929502899197778761945615660047134.9249638941115606306389354461440563212716929502899197778761945615660047 $1.13\times 10^{-378}$
$80$ 135.8744976900983831529154942879362038033124413255633587171958644086561842135.8744976900983831529154942879362038033124413255633587171958644086561842 $8.03\times 10^{-394}$
$81$ 137.2010834668406301716830996653811071822066893852372048164365956206970149137.2010834668406301716830996653811071822066893852372048164365956206970149 $3.01\times 10^{-379}$
$82$ 138.0356990313049020042749619065482729585857816235533229199603352798733897138.0356990313049020042749619065482729585857816235533229199603352798733897 $2.52\times 10^{-379}$
$83$ 139.4105995951457328667296699990735809900833487908551282462172352268054339139.4105995951457328667296699990735809900833487908551282462172352268054339 $1.40\times 10^{-375}$
$84$ 141.9462405516049128078346136764138632079930804595332431920642151809404665141.9462405516049128078346136764138632079930804595332431920642151809404665 $1.77\times 10^{-375}$
$85$ 142.5245124329877516499138087023775533114539886176690581797696212087823187142.5245124329877516499138087023775533114539886176690581797696212087823187 $1.05\times 10^{-380}$
$86$ 143.9615911232873249679700260946989275500715387277893452624660630780480223143.9615911232873249679700260946989275500715387277893452624660630780480223 $1.42\times 10^{-378}$
$87$ 144.9483276132617096716684583857500230575404813669059877812679180958528794144.9483276132617096716684583857500230575404813669059877812679180958528794 $1.37\times 10^{-375}$
$88$ 146.3848096091094233623717818286364786730892073299368170400054147265509705146.3848096091094233623717818286364786730892073299368170400054147265509705 $2.53\times 10^{-378}$
$89$ 147.8797291031824647001555355926284591276893974747116792549589248487923432147.8797291031824647001555355926284591276893974747116792549589248487923432 $4.67\times 10^{-381}$
$90$ 148.6791376989272828780655871531489156030918416831048594377087308271848378148.6791376989272828780655871531489156030918416831048594377087308271848378 $1.26\times 10^{-377}$
$91$ 150.5625106711582684678831224180073436626247034885508876594372802953139708150.5625106711582684678831224180073436626247034885508876594372802953139708 $3.19\times 10^{-380}$
$92$ 152.1147485501154030930412343288739204270768339963739180340167096854953029152.1147485501154030930412343288739204270768339963739180340167096854953029 $5.10\times 10^{-358}$
$93$ 153.2874104994293314602247720515393224042042637960724660197202500041048058153.2874104994293314602247720515393224042042637960724660197202500041048058 $1.82\times 10^{-381}$
$94$ 154.4907156955163636384604753696533838899948856560427610564143071072631799154.4907156955163636384604753696533838899948856560427610564143071072631799 $8.89\times 10^{-382}$
$95$ 155.2420799990558198969997928035576763043815202860039922762556596679189052155.2420799990558198969997928035576763043815202860039922762556596679189052 $5.80\times 10^{-387}$
$96$ 156.4997285439414899913040309653704544908218626069145493265791039370205636156.4997285439414899913040309653704544908218626069145493265791039370205636 $2.98\times 10^{-384}$
$97$ 158.4000270418246442479812803196489920645070375885184935734965513382773251158.4000270418246442479812803196489920645070375885184935734965513382773251 $2.31\times 10^{-375}$
$98$ 159.8727775733136235916063341758884923622323802914269809997571708512468331159.8727775733136235916063341758884923622323802914269809997571708512468331 $1.82\times 10^{-383}$
$99$ 161.2462175616940906041660713748861485716277484845435577262869276737895404161.2462175616940906041660713748861485716277484845435577262869276737895404 $1.88\times 10^{-377}$
$100$ 161.7676625972346591307869118758945874585239556380549233208831548174780607161.7676625972346591307869118758945874585239556380549233208831548174780607 $1.80\times 10^{-377}$
$101$ 163.7576737283108406276969780794254368523306129767593543960804767107794298163.7576737283108406276969780794254368523306129767593543960804767107794298 $3.17\times 10^{-381}$
$102$ 164.6120093931756454229907543346819925745086105838034366833937267160262192164.6120093931756454229907543346819925745086105838034366833937267160262192 $8.98\times 10^{-383}$
$103$ 165.7849132655712156658656369618645110416238945585780065340278032733366523165.7849132655712156658656369618645110416238945585780065340278032733366523 $1.38\times 10^{-370}$
[…]

### 12 Mar 2026 v3

@@ Appendix B Zeros of $L(s,\chi_{5})$ to 70 Decimal Places (Conductor 5, $d=5$)
11 6.64845334472771471612327845997931784729858542324449837237471294671536946.6484533447277147161232784599793178472985854232444983723747129467153694 $1.56\times 10^{-450}$
22 9.83144443288666961634832134745844382188813289401550743042668652269440369.8314444328866696163483213474584438218881328940155074304266865226944036 $2.64\times 10^{-450}$
33 11.958845626083514530265658688262841817293112765723287717381364444659796711.9588456260835145302656586882628418172931127657232877173813644446597967 $2.75\times 10^{-450}$
44 16.033821128384235674593253782248083508749303711393763454343960473652264116.0338211283842356745932537822480835087493037113937634543439604736522641 $4.47\times 10^{-450}$
55 17.566994292325555202701595268144486403485240895508468763100055336820239617.5669942923255552027015952681444864034852408955084687631000553368202396 $3.33\times 10^{-450}$
66 19.540732622784750250378690022998553485526493445007156424107667950060473719.5407326227847502503786900229985534855264934450071564241076679500604737 $5.06\times 10^{-450}$
77 22.227405454459410911877624963081473365280642640560187108353392333259926322.2274054544594109118776249630814733652806426405601871083533923332599263 $6.81\times 10^{-450}$
88 24.588466217408195207656269976081979066247857188503256974763128082376573824.5884662174081952076562699760819790662478571885032569747631280823765738 $6.64\times 10^{-450}$
99 26.776095948004140116523574965270948586289370096100720747141444091727181726.7760959480041401165235749652709485862893700961007207471414440917271817 $7.70\times 10^{-450}$
$10$ 28.461035100177522475186978272325256271200734674528650404495427630566281928.4610351001775224751869782723252562712007346745286504044954276305662819 $6.50\times 10^{-450}$
$11$ 29.707909350480965569230986518657313605970196432418479892145402797880542529.7079093504809655692309865186573136059701964324184798921454027978805425 $7.54\times 10^{-450}$
$12$ 33.000456006870514367949759177211937317376500804631238343447920775517672933.0004560068705143679497591772119373173765008046312383434479207755176729 $1.16\times 10^{-449}$
$13$ 34.728812978904808674143729833981379136564007992078974829862022029197196934.7288129789048086741437298339813791365640079920789748298620220291971969 $1.14\times 10^{-449}$
$14$ 35.868638371812274594595048638876426333384181725562636611431705494857515935.8686383718122745945950486388764263333841817255626366114317054948575159 $8.20\times 10^{-450}$
$15$ 38.129184721436531850151418270372918926216187170437541987827769860251830438.1291847214365318501514182703729189262161871704375419878277698602518304 $1.02\times 10^{-449}$
$16$ 39.560572946403181705055095059954798425566825204646381762421875276690628939.5605729464031817050550950599547984255668252046463817624218752766906289 $1.03\times 10^{-449}$
$17$ 41.842438545791694308509306885312933146785226936948497507540546923147676641.8424385457916943085093068853129331467852269369484975075405469231476766 $1.07\times 10^{-449}$
$18$ 44.031290061441695044700908058424922514756441978971928142545909588119921544.0312900614416950447009080584249225147564419789719281425459095881199215 $1.26\times 10^{-449}$
1919 45.427300082782289388841561909609720649965874840453313174253112295195133545.4273000827822893888415619096097206499658748404533131742531122951951335 $1.28\times 10^{-449}$
2020 46.492727159491405345339199351675833673263826528873731752681947727399791446.4927271594914053453391993516758336732638265288737317526819477273997914 $1.03\times 10^{-449}$
2121 48.345661821067846176548201304493486929945215885626654705310956748852704448.3456618210678461765482013044934869299452158856266547053109567488527044 $1.67\times 10^{-449}$
$22$ 51.087751926746491355258257204131245513931822764598874889643028952539913351.0877519267464913552582572041312455139318227645988748896430289525399133 $1.33\times 10^{-449}$
$23$ 52.125902231316974188604130695930300524969293830813142026749809937021618752.1259022313169741886041306959303005249692938308131420267498099370216187 $1.56\times 10^{-449}$
$24$ 53.830445195442163351054269022338568937339170713109887047091903708283035253.8304451954421633510542690223385689373391707131098870470919037082830352 $1.49\times 10^{-449}$
$25$ 55.589280335404810150964588998492970666633756779445931871191432982474752955.5892803354048101509645889984929706666337567794459318711914329824747529 $1.25\times 10^{-449}$
$26$ 56.838865942890988462381701087979625873787502823421640673322572691726701056.8388659428909884623817010879796258737875028234216406733225726917267010 $1.56\times 10^{-449}$
$27$ 58.386117485358380826793925742317062296785252761744412031426483942467746058.3861174853583808267939257423170622967852527617444120314264839424677460 $1.30\times 10^{-449}$
$28$ 61.138864751921745109095463915572425456187565195524483622740465340768912361.1388647519217451090954639155724254561875651955244836227404653407689123 $1.39\times 10^{-449}$
$29$ 62.132904720233877934071529440383543136671977562274208379683708348891512562.1329047202338779340715294403835431366719775622742083796837083488915125 $1.56\times 10^{-449}$
$30$ 63.709543060981146771798494836335680246320729913038089464768398907875187863.7095430609811467717984948363356802463207299130380894647683989078751878 $1.65\times 10^{-449}$
$31$ 64.637359665374574101088589957514469169614418340871541926806467731667825864.6373596653745741010885899575144691696144183408715419268064677316678258 $1.72\times 10^{-449}$
$32$ 66.747641845937807830433496332522277799056278904231710999500592219175209366.7476418459378078304334963325222777990562789042317109995005922191752093 $2.00\times 10^{-449}$
$33$ 68.590727924825006540833515068030970994309372068387999238433409618374964168.5907279248250065408335150680309709943093720683879992384334096183749641 $3.04\times 10^{-449}$
$34$ 70.115837159193504702254162003288501122534880816675087043045923668594310970.1158371591935047022541620032885011225348808166750870430459236685943109 $1.39\times 10^{-449}$
$35$ 71.605959479243017199477227381292487650674837775870501191061771430349257971.6059594792430171994772273812924876506748377758705011910617714303492579 $2.60\times 10^{-449}$
$36$ 73.350105431468259324986481309687463592380080206128610550028406269746719773.3501054314682593249864813096874635923800802061286105500284062697467197 $1.99\times 10^{-449}$
$37$ 74.293688438016568417563919573323746752656162649388588394478627746168005574.2936884380165684175639195733237467526561626493885883944786277461680055 $1.97\times 10^{-449}$
$38$ 75.534904117228113405345989547687724838297800656054992950178626192713539375.5349041172281134053459895476877248382978006560549929501786261927135393 $2.96\times 10^{-449}$
$39$ 78.072170864382337194051431823672802856073274377354409700849470519410328278.0721708643823371940514318236728028560732743773544097008494705194103282 $1.74\times 10^{-449}$
$40$ 79.786062490097939106371571243171617631261869341105718896147463944735173879.7860624900979391063715712431716176312618693411057188961474639447351738 $2.09\times 10^{-449}$
$41$ 80.567300856410698261373757677661867641972294500859119722736169220177608380.5673008564106982613737576776618676419722945008591197227361692201776083 $1.79\times 10^{-449}$
$42$ 81.923380893154115311967989553905914195519790752160387751303590547918086981.9233808931541153119679895539059141955197907521603877513035905479180869 $1.94\times 10^{-449}$
$43$ 83.699157001874113162883288392240055676480712410299839420963639633825894883.6991570018741131628832883922400556764807124102998394209636396338258948 $2.00\times 10^{-449}$
$44$ 84.952507103754973661846904660485113866224814086703961086051329413504388184.9525071037549736618469046604851138662248140867039610860513294135043881 $2.62\times 10^{-449}$
$45$ 86.809996679338770496962453824123865539107306122702912569153552151872514886.8099966793387704969624538241238655391073061227029125691535521518725148 $3.37\times 10^{-449}$
$46$ 88.281776599709835593621908069365568911371943199556447973711012167487531688.2817765997098355936219080693655689113719431995564479737110121674875316 $3.67\times 10^{-449}$
$47$ 90.376573182128156802963266298552786636345751535570441691904969596990152990.3765731821281568029632662985527866363457515355704416919049695969901529 $2.35\times 10^{-449}$
$48$ 90.715525933146974977269052021613460765711009835158005422179562303740599990.7155259331469749772690520216134607657110098351580054221795623037405999 $1.58\times 10^{-449}$
$49$ 92.450243253774403222656651609940166532625651637342691786476584587897979992.4502432537744032226566516099401665326256516373426917864765845878979799 $2.09\times 10^{-449}$
$50$ 93.394423937828017400982646092448145456052891559124280035714176687016475793.3944239378280174009826460924481454560528915591242800357141766870164757 $2.41\times 10^{-449}$
$51$ 95.970001646037706939832706231680914850344362198662140910938755275316429995.9700016460377069398327062316809148503443621986621409109387552753164299 $2.93\times 10^{-449}$
$52$ 97.331588252898071590582881239627993663344456879633576511321148041911903797.3315882528980715905828812396279936633444568796335765113211480419119037 $2.17\times 10^{-449}$
$53$ 98.227457869901077077061547196562023327793547734165391303478765895425843798.2274578699010770770615471965620233277935477341653913034787658954258437 $2.23\times 10^{-449}$
$54$ 99.838050807536549892458178740443727323094718661873207723162307044585847999.8380508075365498924581787404437273230947186618732077231623070445858479 $2.81\times 10^{-449}$
$55$ 101.1877170969186106879893673648615779856303956787129404948013652775331402101.1877170969186106879893673648615779856303956787129404948013652775331402 $3.14\times 10^{-449}$
$56$ 102.6052879410500105919016338097244199966334774046924761659561257554903097102.6052879410500105919016338097244199966334774046924761659561257554903097 $2.79\times 10^{-449}$
$57$ 103.6661310797677629747891494548931064951051111049995242765910102899167622103.6661310797677629747891494548931064951051111049995242765910102899167622 $3.37\times 10^{-449}$
$58$ 105.9731937977587861297183750394317655839089623673242427608883134019122135105.9731937977587861297183750394317655839089623673242427608883134019122135 $4.22\times 10^{-449}$
$59$ 107.2971036375259575240109457236330672603545936270064737698887145263170400107.2971036375259575240109457236330672603545936270064737698887145263170400 $2.57\times 10^{-449}$
$60$ 108.4981066242615983601377772869913646076568520675142124199793643754286052108.4981066242615983601377772869913646076568520675142124199793643754286052 $2.75\times 10^{-449}$
$61$ 109.5502586071841272439535415119528096196195061162254833594241784691252750109.5502586071841272439535415119528096196195061162254833594241784691252750 $2.77\times 10^{-449}$
$62$ 110.6361976726278518372473829197474157598718943493419745206490444961344615110.6361976726278518372473829197474157598718943493419745206490444961344615 $2.68\times 10^{-449}$
$63$ 112.6157699450494525603420477631256573284904001732412552640056168054705386112.6157699450494525603420477631256573284904001732412552640056168054705386 $3.65\times 10^{-449}$
$64$ 114.2306398723819524205236259115864715634752773232857784617817436415063755114.2306398723819524205236259115864715634752773232857784617817436415063755 $3.51\times 10^{-449}$
$65$ 115.5784139823770731492864674907752976644587354074111098082626495892785778115.5784139823770731492864674907752976644587354074111098082626495892785778 $3.16\times 10^{-449}$
$66$ 116.6356717597837992860716616244380290420850400876742741600872550143161218116.6356717597837992860716616244380290420850400876742741600872550143161218 $2.90\times 10^{-449}$
$67$ 118.3711623063552621519304519550474639914856549056003421824775094839317423118.3711623063552621519304519550474639914856549056003421824775094839317423 $3.74\times 10^{-449}$
$68$ 119.5614916258800030014640284654174824520291754354058843504968760508448926119.5614916258800030014640284654174824520291754354058843504968760508448926 $3.27\times 10^{-449}$
$69$ 120.2817341918527579929633987893433395906991345403172727276829320830538689120.2817341918527579929633987893433395906991345403172727276829320830538689 $2.53\times 10^{-449}$
$70$ 122.0616322838941436366460069970312459167329049609093140518854982869671604122.0616322838941436366460069970312459167329049609093140518854982869671604 $4.20\times 10^{-449}$
$71$ 124.2428371489503574456280462431539381366392074620427334384975693951880955124.2428371489503574456280462431539381366392074620427334384975693951880955 $3.35\times 10^{-449}$
$72$ 125.2815361690544786849959629329922579004633315004421413848554281307878485125.2815361690544786849959629329922579004633315004421413848554281307878485 $4.66\times 10^{-449}$
$73$ 126.3986993784245731137991303757675004132192700408244704238571151980971243126.3986993784245731137991303757675004132192700408244704238571151980971243 $2.92\times 10^{-449}$
$74$ 127.3206348281152351166040431083884153111288443677405189033098037679391083127.3206348281152351166040431083884153111288443677405189033098037679391083 $3.55\times 10^{-449}$
$75$ 129.0483000943208951329053997018161493807340037666172892353902573266257762129.0483000943208951329053997018161493807340037666172892353902573266257762 $3.10\times 10^{-449}$
$76$ 130.1970382668577478738346244894136617196218666973159138428989678080490833130.1970382668577478738346244894136617196218666973159138428989678080490833 $2.86\times 10^{-449}$
$77$ 132.0860053031175217593818811988220993124226919197273881005207706354373250132.0860053031175217593818811988220993124226919197273881005207706354373250 $2.98\times 10^{-449}$
$78$ 133.0302283883319444007316023118379973226202504376344453082369748548055879133.0302283883319444007316023118379973226202504376344453082369748548055879 $4.95\times 10^{-449}$
$79$ 134.9249638941115606306389354461440563212716929502899197778761945615660048134.9249638941115606306389354461440563212716929502899197778761945615660048 $4.58\times 10^{-449}$
$80$ 135.8744976900983831529154942879362038033124413255633587171958644086561842135.8744976900983831529154942879362038033124413255633587171958644086561842 $4.13\times 10^{-449}$
$81$ 137.2010834668406301716830996653811071822066893852372048164365956206970149137.2010834668406301716830996653811071822066893852372048164365956206970149 $4.07\times 10^{-449}$
$82$ 138.0356990313049020042749619065482729585857816235533229199603352798733898138.0356990313049020042749619065482729585857816235533229199603352798733898 $4.70\times 10^{-449}$
$83$ 139.4105995951457328667296699990735809900833487908551282462172352268054340139.4105995951457328667296699990735809900833487908551282462172352268054340 $4.09\times 10^{-449}$
$84$ 141.9462405516049128078346136764138632079930804595332431920642151809404665141.9462405516049128078346136764138632079930804595332431920642151809404665 $3.57\times 10^{-449}$
$85$ 142.5245124329877516499138087023775533114539886176690581797696212087823188142.5245124329877516499138087023775533114539886176690581797696212087823188 $3.43\times 10^{-449}$
$86$ 143.9615911232873249679700260946989275500715387277893452624660630780480223143.9615911232873249679700260946989275500715387277893452624660630780480223 $4.72\times 10^{-449}$
$87$ 144.9483276132617096716684583857500230575404813669059877812679180958528794144.9483276132617096716684583857500230575404813669059877812679180958528794 $4.18\times 10^{-449}$
$88$ 146.3848096091094233623717818286364786730892073299368170400054147265509706146.3848096091094233623717818286364786730892073299368170400054147265509706 $3.81\times 10^{-449}$
$89$ 147.8797291031824647001555355926284591276893974747116792549589248487923432147.8797291031824647001555355926284591276893974747116792549589248487923432 $5.20\times 10^{-449}$
$90$ 148.6791376989272828780655871531489156030918416831048594377087308271848378148.6791376989272828780655871531489156030918416831048594377087308271848378 $5.09\times 10^{-449}$
$91$ 150.5625106711582684678831224180073436626247034885508876594372802953139708150.5625106711582684678831224180073436626247034885508876594372802953139708 $2.95\times 10^{-428}$
$92$ 152.1147485501154030930412343288739204270768339963739180340167096854953029152.1147485501154030930412343288739204270768339963739180340167096854953029 $6.49\times 10^{-577}$
$93$ 153.2874104994293314602247720515393224042042637960724660197202500041048059153.2874104994293314602247720515393224042042637960724660197202500041048059 $4.83\times 10^{-449}$
$94$ 154.4907156955163636384604753696533838899948856560427610564143071072631800154.4907156955163636384604753696533838899948856560427610564143071072631800 $3.47\times 10^{-449}$
$95$ 155.2420799990558198969997928035576763043815202860039922762556596679189052155.2420799990558198969997928035576763043815202860039922762556596679189052 $4.14\times 10^{-449}$
$96$ 156.4997285439414899913040309653704544908218626069145493265791039370205637156.4997285439414899913040309653704544908218626069145493265791039370205637 $6.47\times 10^{-449}$
$97$ 158.4000270418246442479812803196489920645070375885184935734965513382773252158.4000270418246442479812803196489920645070375885184935734965513382773252 $9.41\times 10^{-449}$
$98$ 159.8727775733136235916063341758884923622323802914269809997571708512468331159.8727775733136235916063341758884923622323802914269809997571708512468331 $6.21\times 10^{-449}$
$99$ 161.2462175616940906041660713748861485716277484845435577262869276737895404161.2462175616940906041660713748861485716277484845435577262869276737895404 $3.23\times 10^{-449}$
$100$ 161.7676625972346591307869118758945874585239556380549233208831548174780607161.7676625972346591307869118758945874585239556380549233208831548174780607 $4.89\times 10^{-449}$
$101$ 163.7576737283108406276969780794254368523306129767593543960804767107794299163.7576737283108406276969780794254368523306129767593543960804767107794299 $5.58\times 10^{-449}$
$102$ 164.6120093931756454229907543346819925745086105838034366833937267160262192164.6120093931756454229907543346819925745086105838034366833937267160262192 $4.83\times 10^{-449}$
$103$ 165.7849132655712156658656369618645110416238945585780065340278032733366524165.7849132655712156658656369618645110416238945585780065340278032733366524 $5.17\times 10^{-449}$
[…]